OpenStudy (anonymous):

What is the quotient in simplified form? State any restrictions on the variable. x^2-16/x^3+5x+6 divided by x^3+5x+4/x^2-2x-8

OpenStudy (anonymous):

\[\frac{ x^{2}-16 }{ x^{2}+5x+6 } \div \frac{ a^{2}+5x+4 }{ x^{2}-2x-8 }\]

OpenStudy (anonymous):

@johnweldon1993 @jim_thompson5910

OpenStudy (anonymous):

invert and multiply, but really factor and cancel

OpenStudy (anonymous):

can you walk me through it??

OpenStudy (anonymous):

\[\frac{ x^{2}-16 }{ x^{2}+5x+6 } \div \frac{ a^{2}+5x+4 }{ x^{2}-2x-8 }\] \[=\frac{ x^{2}-16 }{ x^{2}+5x+6 } \times \frac{x^{2}-2x-8 }{ x^{2}+5x+4 }\]

OpenStudy (anonymous):

i assume the \(a\) was a typo, and it should be an \(x\) right?

OpenStudy (anonymous):

now this problem was cooked so you don't actually multiply anything factor each binomial , then cancel common factors

OpenStudy (anonymous):

\[=\frac{ x^{2}-16 }{ x^{2}+5x+6 } \times \frac{x^{2}-2x-8 }{ x^{2}+5x+4 }\] \[=\frac{(x+4)(x-4)(x-6)(x+2)}{(x+3)(x+2)(x+4)(x+1)}\] all factored out

OpenStudy (anonymous):

ach i made a mistake!!

OpenStudy (anonymous):

\[\large =\frac{(x+4)(x-4)(x-4)(x+2)}{(x+3)(x+2)(x+4)(x+1)}\] that's better

OpenStudy (anonymous):

okay hold on. let me look at this and see if i can catch on.&&yes that z was supposed to be an x

OpenStudy (anonymous):

ok take your time step one was to flip the second one step two was to factor each quadratic the last step, which we didn't get to yet, is to cancel common factors

OpenStudy (anonymous):

\[\frac{ (x-4)(x-4) }{ (x+3)(x+1) }\]

OpenStudy (anonymous):

is this what i should get after canceling out common factors?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and finally you need all the "restrictions" find the zeros of the original denominator \[(x+3)(x+2)(x+4)(x+1)\] and those are the restricted values i.e. \[-3,-2,-1,-4\]

OpenStudy (anonymous):

would 4 be one?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

there is no \(x-4\) in the denominator